OpenStudy (anonymous):

a math test is made up of 5point and 10point questions. there are 24 questions on the test. the total number of possible points is 150points What is the total value of all the 10-point questions on test

OpenStudy (anonymous):

let x be the no. of 5 points question and y be the no. of 10points question. x + y = 24 , x = 24-y-----------> i.) 5x + 10y = 150 5(24-y) +10y=150 120-5y+10y=150 5y=150-120 5y=30 y=6 Total value of 6 10point question = 6*10 = 60

OpenStudy (anonymous):

thanks